On a windy day, a boat traveled 210 km in 15 hours. When travelling against the wind, it took 21hours for the boat to travel back. What was the rate of the boat in still air and what was the velocity of the wind that hindered the speed of the boat?
Found 2 solutions by ankor@dixie-net.com, ikleyn:
Answer by ankor@dixie-net.com(22740)

On a windy day, a boat traveled 210 km in 15 hours.
When travelling against the wind, it took 21 hours for the boat to travel back.
What was the rate of the boat in still air and what was the velocity of the wind that hindered the speed of the boat?
:
let s = the normal speed of the boat on a windless day
let w = the rate of the wind
then
(s+w) = the effective speed with the wind
and
(s-w) = the effective speed against the wind
:
Write distance equation for each way. dist = time * speed
15(s+w) = 210
21(s-w) = 210
simplify both equations, divide the first by 15, divide the second by 21
s + w = 14
s - w = 10
-------------addition eliminates w find s
2s + 0 = 24
s = 24/2
s = 12 km/h is the speed of the boat
:
Find the rate of the wind using the 1st simplified equation
12 + w = 14
w = 14 - 12
w = 2 km/hr is the rate of the wind
:
:
Check this in the statement
" On a windy day, a boat traveled 210 km in 15 hours."
15(12 + 2) =
15(14) = 210 km

Answer by ikleyn(52799)

.
On a windy day, a boat traveled 210 km in 15 hours. When travelling against the wind, it took 21hours for the boat to travel back.
What was the rate of the boat in still air and what was the velocity of the wind that hindered the speed of the boat?
~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~


From the Physics point of view, this problem is   ABSOLUTE NONSENSE.


I am repeating:     From the Physics point of view, this problem is   ABSOLUTE NONSENSE.


        I am repeating it again:         From the Physics point of view, this problem is   ABSOLUTE NONSENSE.



The solution by the tutor  "ankor@dixie-net.com"  is   ABSOLUTE NONSENSE,  too.  Unfortunately.



The error and the reason for my comment is this:


    If the speed of the boat in still water is "s" and the wind speed is "w" then the effective speed of the boat  "with the wind"  IS  NOT  (s+w).


    If the speed of the boat in still water is "s" and the wind speed is "w" then the effective speed of the boat  "against the wind"  IS  NOT  (s-w).


It is so obvious that does not deserve any further discussions.


This problem  //or its close version//  just came to the forum once before,
and I reacted with the same comments.


I warned the author that any further mentions or the trials to create a discussion I will cancel immediately.


So stupid it is.


Now it came again.


My warning still remains in force.   And will remain forever.


By entering with such post,  you,  the author,  demonstrate and INSIST  that your knowledge of  Physics is  ZERO,  if not NEGATIVE.